# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v1

@@ D-F Monotonicity and Convexity of SP Cost
D-G Proof Sketch of Theorem 1
@@
  • By Lemma 2, the congestion contributions to $\Delta U_{\mathbbm{j}}$ and $\Delta\Phi$ have the same sign.
  • By Appx. D-F, the SP-cost variation responds monotonically and convexly to the direction of load shift, further aligning the sign of $\Delta\Phi$ with that of $\Delta U_{\mathbbm{j}}$.
  • Payment differences $(p_{\mathbbm{j},\mathbbm{i}^{\prime}}-p_{\mathbbm{j},\mathbbm{i}})$ only redistribute welfare and are excluded from $\Delta\Phi$ (Appx. D-I), so they do not affect the ordinal comparison.
Hence, $\Delta U_{\mathbbm{j}}>0$ implies $\Delta\Phi>0$, which proves that $\Phi(\bm{\pi})$ is an ordinal potential for the considered game, as stated in Theorem 1.
D-H Finite-Improvement Property and Existence of NE
@@
Let $\mathcal{A}_{\mathbbm{j}}=\{0\}\cup\mathcal{S}^{\mathrm{On}}$ be SD $u^{\mathrm{On}}_{\mathbbm{j}}$’s finite action set, and $\mathcal{A}=\prod_{\mathbbm{j}}\mathcal{A}_{\mathbbm{j}}$ the joint action space. Consider any strict better-response dynamics in which, at each step, some SD switches to an action that strictly improves its utility. By Theorem 1, every such deviation strictly increases the potential $\Phi(\bm{\pi})$.
Since $\mathcal{A}$ is finite, the sequence of potential values cannot increase indefinitely. The process must terminate after finitely many steps at a profile $\bm{\pi}^{\star}$ where no SD can further increase its utility via a unilateral deviation. By definition, $\bm{\pi}^{\star}$ is a pure-strategy NE. Moreover, any local maximizer of $\Phi(\bm{\pi})$ is a pure-strategy NE, because any unilateral move would reduce (or not increase) $\Phi$, and hence would not increase the deviator’s utility in an ordinal-potential game.
This finite-improvement property underpins PG-BRD: as long as SDs occasionally update their actions based on (approximate) best responses, the algorithm converges in finitely many steps to a stable congestion-aware offloading configuration.
D-I Treatment of Monetary Payments in the Potential Function

### v2

@@ E-F Monotonicity and Convexity of SP Cost
E-G Proof Sketch of Proposition 7
@@
  • By Proposition 9, the congestion contributions to $\Delta U_{\mathbbm{j}}$ and $\Delta\Phi$ have the same sign.
  • By Appx. E-F, the SP-cost variation responds monotonically and convexly to the direction of load shift, further aligning the sign of $\Delta\Phi$ with that of $\Delta U_{\mathbbm{j}}$.
  • Payment differences $(p_{\mathbbm{j},\mathbbm{i}^{\prime}}-p_{\mathbbm{j},\mathbbm{i}})$ only redistribute welfare and are excluded from $\Delta\Phi$ (Appx. E-I), so they do not affect the ordinal comparison.
Hence, $\Delta U_{\mathbbm{j}}>0$ implies $\Delta\Phi>0$, which proves that $\Phi(\bm{\pi})$ is an ordinal potential for the considered game, as stated in Proposition 7.
E-H Finite-Improvement Property and Existence of NE
@@
Let $\mathcal{A}_{\mathbbm{j}}=\{0\}\cup\mathcal{S}^{\mathrm{On}}$ be SD $u^{\mathrm{On}}_{\mathbbm{j}}$’s finite action set, and $\mathcal{A}=\prod_{\mathbbm{j}}\mathcal{A}_{\mathbbm{j}}$ the joint action space. Consider any strict better-response dynamics in which, at each step, some SD switches to an action that strictly improves its utility. By Proposition 7, every such deviation strictly increases the potential $\Phi(\bm{\pi})$.
Since $\mathcal{A}$ is finite, the sequence of potential values cannot increase indefinitely. The process must terminate after finitely many steps at a profile $\bm{\pi}^{\star}$ where no SD can further increase its utility via a unilateral deviation. By definition, $\bm{\pi}^{\star}$ is a pure-strategy NE. Moreover, any local maximizer of $\Phi(\bm{\pi})$ is a pure-strategy NE, because any unilateral move would reduce (or not increase) $\Phi$, and hence would not increase the deviator’s utility in an ordinal-PG.
This finite-improvement property underpins PG-BRD: as long as SDs occasionally update their actions based on best responses, the algorithm converges in finitely many steps to a stable congestion-aware offloading configuration.
E-I Treatment of Monetary Payments in the Potential Function